The question says find the x-intercepts y=4x^2+5x-6
y=4x^2+5x-6
0=4X^2+5X-6
(4x^2+-3x)+(8x+-6)
x(4x+-3)+2(4x+-3)
(x+2) (4x+-3)
Ok I can get to here but I need to solve for the X-intercepts and I'm stuck
Please help

Answer by uma(370)

You are absolutely right.
Now equate it to 0.
==> (x+2)(4x-3) = 0
The product cannot be 0 unless either of the factors is 0.
==> x + 2 = 0 or 4x - 3 = 0
==> x = -2 or 4x = 3
==> x = -2 or x = 3/4
Thus the x intercepts are -2 and 3/4 or 0.75

Good Luck!!!